# Products of two orthogonal projections

### v1

@@ 6. Model projections
Recall that the model space $\mathcal{Q}_{\varphi}$ corresponding to an inner function $\varphi\in H^{\infty}(\mathbb{D}^{n})$ is defined by
@@
We refer to $P_{\mathcal{Q}_{\varphi}}$ the projection onto the model space $\mathcal{Q}_{\varphi}$ as a model projection. This section seeks to understand operators that can be represented as the product of two model projections. Similarly to inner projections, if a nonzero operator $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$ can be represented as a product of two model projections, we can conclude that both the $\text{ran}T$ and $\text{ran}T^{*}$ are nonzero. For each nonzero $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$, we define
@@
Clearly, the set $\mathcal{J}_{T}$ mentioned above could potentially be empty (for instance, whenever $T$ is a co-isometry). However, if the set is not trivial, then, as shown in the proof of Lemma 5.2, one may conclude that the gcd of $\mathcal{J}_{T}$ exists.
@@
Similar to the proof of Lemma 5.2, we also have the following existential result. However, in this case, it is necessary to assume that the set $\mathcal{J}_{T}$ is nonempty.
@@
We are now prepared to present the result related to the products of the model operators. The proof most often deviates from that of Theorem 5.3. Also note that $\psi_{T^{*}}$ is the gcd of $\mathcal{J}_{T^{*}}$.
@@
Suppose there are inner functions $\varphi_{1},\varphi_{2}\in H^{\infty}(\mathbb{D}^{n})$ such that $T=P_{\mathcal{Q}_{\varphi_{1}}}P_{\mathcal{Q}_{\psi_{\varphi_{2}}}}$. In this case, $\varphi_{1}\in\mathcal{J}_{T}$ and $\varphi_{2}\in\mathcal{J}_{T^{*}}$ together imply that $\mathcal{J}_{T}$ and $\mathcal{J}_{T^{*}}$ are nonempty sets, and hence $\psi_{T}$ and $\psi_{T^{*}}$ are nonzero inner functions in $H^{\infty}(\mathbb{D}^{n})$. Since $\psi_{T}$ in $\mathcal{J}_{T}$, and an inner function, it follows that $\psi_{T}H^{2}(\mathbb{D}^{n})\subseteq\ker T^{*}$, and hence
@@
that is, $\text{ran}T\subseteq\mathcal{Q}_{\psi_{T}}$. Now, we know in general that $\text{ran}TP_{\mathcal{Q}_{T^{*}}}\subseteq\text{ran}T$, which immediately implies that $\text{ran}TP_{\mathcal{Q}_{\psi_{T^{*}}}}\subseteq\mathcal{Q}_{\psi_{T}}$, and hence we have the identity
@@
Next, as $\psi_{T^{*}}\in\mathcal{J}_{T^{*}}$, we have $\psi_{T^{*}}H^{2}(\mathbb{D}^{n})\subseteq\ker T$, equivalently, $(\ker T)^{\perp}\subseteq\mathcal{Q}_{\psi_{T^{*}}}$. This implies $P_{\mathcal{Q}_{\psi_{T^{*}}}}P_{(\ker T)^{\perp}}=P_{(\ker T)^{\perp}}$, and hence $TP_{\mathcal{Q}_{\psi_{T^{*}}}}=TP_{(\ker T)^{\perp}}=T$. This combined with $P_{\mathcal{Q}_{\psi_{T}}}TP_{\mathcal{Q}_{\psi_{T^{*}}}}=TP_{\mathcal{Q}_{\psi_{T^{*}}}}$ yields
@@
Looking at the product of projections $P_{\mathcal{Q}_{\psi_{T}}}P_{\mathcal{Q}_{\varphi_{1}}}$, we observe that both $\psi_{T}$ and $\varphi_{1}$ are in $\mathcal{J}_{T}$, which implies by the definition of gcd that $\varphi_{1}$ divides $\psi_{T}$; equivalently, $\mathcal{Q}_{\psi_{T}}\subseteq\mathcal{Q}_{\varphi_{1}}$. This gives $P_{\mathcal{Q}_{\psi_{T}}}P_{\mathcal{Q}_{\varphi_{1}}}=P_{\mathcal{Q}_{\psi_{T}}}$, and also, similarly, $P_{\mathcal{Q}_{\psi_{\varphi_{2}}}}P_{\mathcal{Q}_{\psi_{T^{*}}}}=P_{\mathcal{Q}_{\psi_{T^{*}}}}$. This completes the proof of the fact that $T=P_{\mathcal{Q}_{\psi_{T}}}P_{\mathcal{Q}_{\psi_{T^{*}}}}$.
At this time, we will make a general observation regarding kernels of product of two projections: When $T\in\mathcal{B}(\mathcal{H})$ is expressed as a product of projections, $T=P_{1}P_{2}$, we obtain:
@@
Indeed, if $P_{1}P_{2}h=0$, then we write $h=h_{r}\oplus h_{n}\in\text{ran}P_{2}\oplus\ker P_{2}$. Therefore
@@
This, when combined with Corollary 2.5 in the context of inner projections, results in the following characterization of operators as products of two inner projections:
@@
Finally, we again consider model projections. Corollary 2.5 and the identity (6.1) yields:
@@ 7. Blaschke products
is popularly known as a Blaschke factor. The finite Blaschke product is defined by
@@
if and only if $\varphi$ is a finite Blaschke product. The lemma that follows is elementary but of general interest and may have a broader scope when considered in its full generality.
@@
This implies, in particular, that $|\mathcal{Z}(f)|\leq m_{1}-1$. On the other hand, $f\in\varphi_{2}H^{2}(\mathbb{D})$ implies that, counted with multiplicities, $f$ has at least $m_{2}$ zeroes (as $|\mathcal{Z}(\varphi_{2})|\leq|\mathcal{Z}(f)|$). Therefore, we have
@@
and consequently, $m_{2}<m_{1}$; a contradiction. This forces that $f=0$. Now, we turn to prove the converse direction. Let $\mathcal{Q}_{\varphi_{1}}\cap\varphi_{2}H^{2}(\mathbb{D})=\{0\}$, and suppose $m_{1}>m_{2}$. Assume that $\alpha_{1},\ldots,\alpha_{s}$ be distinct zeros of $\varphi_{2}$ with multiplicities $n_{1},\ldots,n_{s}$, respectively. Set
@@
The later identity implies $\ker T=\mathcal{Q}_{\varphi_{T^{*}}}$. We also know that
@@
In other words, we have $\mathcal{Q}_{\varphi_{2}}\cap\varphi_{1}H^{2}(\mathbb{D})=\{0\}$, and hence, Lemma 7.1 implies that $|\mathcal{Z}(\varphi_{2})|\leq|\mathcal{Z}(\varphi_{1})|$. Similarly, as $T^{*}$ is also a product of two inner projections, working as above, we find that $|\mathcal{Z}(\varphi_{1})|\leq|\mathcal{Z}(\varphi_{2})|$. This yields $|\mathcal{Z}(\varphi_{2})|=|\mathcal{Z}(\varphi_{1})|$ – a contradiction. This completes the proof of the proposition.
@@
Suppose $T$ is a product of two inner projections. We again use the identity that
@@
For examples of bounded linear operators on $H^{2}(\mathbb{D})$ that meet the criteria stated in the above result, we note that for every inner function $\theta\in H^{\infty}(\mathbb{D})$, there exists a Hankel operator $H_{\varphi}$, $\varphi\in L^{\infty}(\mathbb{T})$, such that [21, page 15]

### v2

@@ 6. Model projections
Recall that the model space $\mathcal{Q}_{\varphi}$ corresponding to an inner function $\varphi\in H^{\infty}(\mathbb{D}^{n})$ is the quotient space defined by
@@
We refer to $P_{\mathcal{Q}_{\varphi}}$ the projection onto the model space $\mathcal{Q}_{\varphi}$ as a model projection. This section seeks to understand operators that can be represented as products of two model projections. Similarly to inner projections, if a nonzero operator $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$ can be represented as a product of two model projections, we can conclude that both the $\text{ran}T$ and $\text{ran}T^{*}$ are nonzero. For each nonzero $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$, we define
@@
Clearly, the set $\mathcal{J}_{T}$ defined above could potentially be empty (for instance, whenever $T$ is a co-isometry). However, if the set is not trivial, then, as shown in the proof of Lemma 5.2, one may conclude that the gcd of $\mathcal{J}_{T}$ exists.
@@
Similar to the proof of Lemma 5.2, we also have the following existential result. However, in this case, it is necessary to assume that the set $\mathcal{J}_{T}$ is nonempty (note that $\mathcal{I}_{T}$ is automatically nonempty).
@@
We are now prepared to present the result related to the products of model operators. The proof most often deviates from that of Theorem 5.3. Also note that $\psi_{T^{*}}$ is the gcd of $\mathcal{J}_{T^{*}}$ (the adjoint version of $\mathcal{J}_{T}$).
@@
Suppose there are inner functions $\varphi_{1},\varphi_{2}\in H^{\infty}(\mathbb{D}^{n})$ such that $T=P_{\mathcal{Q}_{\varphi_{1}}}P_{\mathcal{Q}_{\psi_{\varphi_{2}}}}$. In this case, $\varphi_{1}\in\mathcal{J}_{T}$ (note that $\mathcal{Q}_{\varphi_{1}}=(\varphi_{1}H^{2}(\mathbb{D}^{n}))^{\perp}$) and $\varphi_{2}\in\mathcal{J}_{T^{*}}$ together imply that $\mathcal{J}_{T}$ and $\mathcal{J}_{T^{*}}$ are nonempty sets, and hence, Lemma 6.2 implies that $\psi_{T}$ and $\psi_{T^{*}}$ are nonzero inner functions in $H^{\infty}(\mathbb{D}^{n})$. Since $\psi_{T}$ in $\mathcal{J}_{T}$, and an inner function, it follows that $\psi_{T}H^{2}(\mathbb{D}^{n})\subseteq\ker T^{*}$, and hence
@@
that is, $\text{ran}T\subseteq\mathcal{Q}_{\psi_{T}}$. Now, we know in general that $\text{ran}(TP_{\mathcal{Q}_{T^{*}}})\subseteq\text{ran}T$, which immediately implies that $\text{ran}(TP_{\mathcal{Q}_{\psi_{T^{*}}}})\subseteq\mathcal{Q}_{\psi_{T}}$, and hence we have the identity
@@
Next, as $\psi_{T^{*}}\in\mathcal{J}_{T^{*}}$, we have $\psi_{T^{*}}H^{2}(\mathbb{D}^{n})\subseteq\ker T$, equivalently, $(\ker T)^{\perp}\subseteq\mathcal{Q}_{\psi_{T^{*}}}$. This implies $P_{\mathcal{Q}_{\psi_{T^{*}}}}P_{(\ker T)^{\perp}}=P_{(\ker T)^{\perp}}$, and hence
$TP_{\mathcal{Q}_{\psi_{T^{*}}}}=TP_{(\ker T)^{\perp}}=T.$
This combined with $P_{\mathcal{Q}_{\psi_{T}}}TP_{\mathcal{Q}_{\psi_{T^{*}}}}=TP_{\mathcal{Q}_{\psi_{T^{*}}}}$ yields
@@
Looking at the product of two projections $P_{\mathcal{Q}_{\psi_{T}}}P_{\mathcal{Q}_{\varphi_{1}}}$, we observe that both $\psi_{T}$ and $\varphi_{1}$ are in $\mathcal{J}_{T}$, which implies by the definition of gcd that $\varphi_{1}$ divides $\psi_{T}$; equivalently, $\mathcal{Q}_{\psi_{T}}\subseteq\mathcal{Q}_{\varphi_{1}}$. This gives $P_{\mathcal{Q}_{\psi_{T}}}P_{\mathcal{Q}_{\varphi_{1}}}=P_{\mathcal{Q}_{\psi_{T}}}$, and also, similarly, $P_{\mathcal{Q}_{\psi_{\varphi_{2}}}}P_{\mathcal{Q}_{\psi_{T^{*}}}}=P_{\mathcal{Q}_{\psi_{T^{*}}}}$. This completes the proof of the fact that $T=P_{\mathcal{Q}_{\psi_{T}}}P_{\mathcal{Q}_{\psi_{T^{*}}}}$.
At this time, we will make a general observation regarding kernels of product of two projections: When $T\in\mathcal{B}(\mathcal{H})$ is expressed as a product of two projections, $T=P_{1}P_{2}$, we obtain:
@@
Indeed, if $P_{1}P_{2}h=0$ for some $h\in\mathcal{H}$, then we write $h=h_{r}\oplus h_{n}\in\text{ran}P_{2}\oplus\ker P_{2}$. Therefore
@@
This, when combined with Proposition 2.4 in the context of inner projections, results in the following characterization of operators as products of two inner projections:
@@
Proof.
If $T$ is a product of two inner projections, then Theorem 5.3 implies
$T=P_{\varphi_{T}H^{2}(\mathbb{D}^{n})}P_{\varphi_{T^{*}}H^{2}(\mathbb{D}^{n})}.$
The representation of $\ker T$ then follows from (6.1). For the other identity, we compute:
$TT^{*}=P_{\varphi_{T}H^{2}(\mathbb{D}^{n})}P_{\varphi_{T^{*}}H^{2}(\mathbb{D}^{n})}P_{\varphi_{T}H^{2}(\mathbb{D}^{n})}=P_{\varphi_{T}H^{2}(\mathbb{D}^{n})}T^{*},$
We now turn to prove the sufficient part. By Proposition 2.4 and $TT^{*}=P_{\varphi_{T}H^{2}(\mathbb{D}^{n})}T^{*}$, we know that
$T=P_{\varphi_{T}H^{2}(\mathbb{D}^{n})}P_{(\ker T)^{\perp}}.$
Define
$\tilde{T}=P_{\varphi_{T}H^{2}(\mathbb{D}^{n})}P_{\varphi_{T^{*}}H^{2}(\mathbb{D}^{n})}.$
We claim that $T=\tilde{T}$. The given assumption about $\ker T$ and the identity (6.1) yield
$\ker T=\ker\tilde{T}.$
In particular, $\mathcal{Q}_{\varphi_{T^{*}}}\subseteq\ker T$, that is, $(\ker T)^{\perp}\subseteq\varphi_{T^{*}}H^{2}(\mathbb{D}^{n})$. This implies $P_{(\ker T)^{\perp}}=P_{\varphi_{T^{*}}H^{2}(\mathbb{D}^{n})}P_{(\ker T)^{\perp}}$, and hence
$T=P_{\varphi_{T}H^{2}(\mathbb{D}^{n})}P_{(\ker T)^{\perp}}=P_{\varphi_{T}H^{2}(\mathbb{D}^{n})}P_{\varphi_{T^{*}}H^{2}(\mathbb{D}^{n})}P_{(\ker T)^{\perp}}=P_{\varphi_{T}H^{2}(\mathbb{D}^{n})}P_{\varphi_{T^{*}}H^{2}(\mathbb{D}^{n})}=\tilde{T}.$
This completes the proof of the theorem.
Finally, we again consider model projections. The proof in this case is similar to the previous one.
@@ 7. Blaschke products
is popularly known as a Blaschke factor. A finite Blaschke product is an inner function $\varphi\in H^{\infty}(\mathbb{D})$ such that
@@
if and only if $\varphi$ is a finite Blaschke product. The lemma that follows is elementary but of general interest and may have a broader scope when considered in its full generality. The cardinality of a set $A$ is denoted by the notation $|A|$.
@@
This implies, in particular, that $|\mathcal{Z}(f)|\leq m_{1}-1$. On the other hand, $f\in\varphi_{2}H^{2}(\mathbb{D})$ implies that, counted with multiplicities, $f$ has at least $m_{2}$ zeros (as $|\mathcal{Z}(\varphi_{2})|\leq|\mathcal{Z}(f)|$). Therefore, we have
@@
and consequently, $m_{2}<m_{1}$; a contradiction. This forces that $f=0$. Now, we turn to prove the converse direction. Let $\mathcal{Q}_{\varphi_{1}}\cap\varphi_{2}H^{2}(\mathbb{D})=\{0\}$, and suppose $m_{1}>m_{2}$. Assume that $\alpha_{1},\ldots,\alpha_{s}$ are distinct zeros of $\varphi_{2}$ with multiplicities $n_{1},\ldots,n_{s}$, respectively. Set
@@
In the following, we relate canonical factorizations of the products of inner projections with corresponding range spaces.
@@
The later identity implies $\ker T=\mathcal{Q}_{\varphi_{T^{*}}}$. By (6.1), we also know that
@@
In other words, we have $\mathcal{Q}_{\varphi_{2}}\cap\varphi_{1}H^{2}(\mathbb{D})=\{0\}$, and hence, Lemma 7.1 implies that $|\mathcal{Z}(\varphi_{2})|\leq|\mathcal{Z}(\varphi_{1})|$. Similarly, as $T^{*}$ is also a product of two inner projections, working as above, we find that $|\mathcal{Z}(\varphi_{1})|\leq|\mathcal{Z}(\varphi_{2})|$. This yields $|\mathcal{Z}(\varphi_{2})|=|\mathcal{Z}(\varphi_{1})|$ – a contradiction, which completes the proof of the proposition.
For operators that are products of two inner projections, the corresponding kernel spaces have a specific relationship with the set of all inner functions.
@@
Suppose $T$ is a product of two inner projections. We have, in particular, that $T=P_{\varphi_{T}H^{2}(\mathbb{D}^{n})}P_{\varphi_{T^{*}}H^{2}(\mathbb{D}^{n})}$. We again use the identity (6.1) and observe that
@@
As a result, we have the following result that says when an operator on $H^{2}(\mathbb{D})$ can’t be represented as a product of two inner projections.
@@
For examples of bounded linear operators on $H^{2}(\mathbb{D})$ that meet the criteria stated in the above result, we note that for every inner function $\theta\in H^{\infty}(\mathbb{D})$, there exists a Hankel operator $H_{\varphi}$, $\varphi\in L^{\infty}(\mathbb{T})$, such that [20, page 15]